OpenStudy (anonymous):

Find an equation of the line having the given slope and containing the given point m=5/6,(7,-1)

OpenStudy (anonymous):

y=5/6x-6.83

jimthompson5910 (jim_thompson5910):

\[y-y_1=m(x-x_1)\] \[y-(-1)=\frac{5}{6}(x-7)\] \[y+1=\frac{5}{6}(x-7)\] \[y+1=\frac{5}{6}x-(7)\frac{5}{6}\] \[y+1=\frac{5}{6}x-\frac{35}{6}\] \[y=\frac{5}{6}x-\frac{35}{6}-1\] \[y=\frac{5}{6}x-\frac{41}{6}\]

OpenStudy (anonymous):

which is y=5/6x-6.83 yours is just in fractions